What is the effective annual rate on an investment that pays an interest of 6.25% continuously?

Select one:

a. 6.45%

b. 6.72%

c. 6.35%

d. 4.47%

e. 8.68%
Answer #1

6.25% continuously compounding:

= e^0.0.0625-1

= 1.064494-1

= 0.064494*100

= 6.4494%   6.45%.

Hence, the answer is option a.

In case of any query, kindly comment on the solution.